OpenStudy (anonymous):

Use basic identities to simplify this expression 1/cot^2x +sec sin and help me like @doulikepiecauseidont did

OpenStudy (anonymous):

\[\Large \frac{1}{\cot^2\theta} +\sec \theta \sin \theta\]??

OpenStudy (anonymous):

Yeah